OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 16x + 64= 0 A. 8 B. 16 C. –8 D. –16

OpenStudy (anonymous):

@Barrelracer011

OpenStudy (barrelracer011):

Well do you know what perfect squares are?

OpenStudy (anonymous):

8

OpenStudy (barrelracer011):

I would believe it to be A

OpenStudy (anonymous):

Thank You

OpenStudy (barrelracer011):

No problem

OpenStudy (anonymous):

X^2+16x+___= -64 16^2=256\2=128 X^2+16x+128=-64+128 (X-8)^2= 64

OpenStudy (barrelracer011):

Yes but then you do the perfect square of 64 and get 8

OpenStudy (barrelracer011):

@reyannagirl

OpenStudy (anonymous):

X=8

OpenStudy (barrelracer011):

Yes so it is A